### v3

@@ 2.2 Metric subregularity of residual mapping
(6) ${\rm dist}(x,\mathcal{S}^{*})\leq\kappa[r(x)]^{\gamma}\quad{\rm for\ all}\ x\in\mathbb{B}(\overline{x},\delta),$
which is precisely the local Hölderian error bound of order $\gamma$ on $\mathcal{S}^{*}$ at $\overline{x}$. Notice that Bai et al. [5] proposed a weaker sufficient condition, i.e. directional pseudo-normality, for the subregularity of constraint system $P(x)\in\Lambda$, where $P\!:\mathbb{X}\to\mathbb{Y}$ is continuous near the point of interest and $\Lambda\subset\mathbb{Y}$ is closed near the point of interest. To apply their results to provide some checkable conditions for the subregularity of system $R(x)=0$, we here recall the tangent, regular normal and (limiting) directional normal cones. For a closed set $S\subset\mathbb{X}$ and a point $x\in S$, the tangent cone to $S$ at $x$ is defined as
$\mathcal{T}_{S}(x):=\big{\{}w\in\mathbb{X}\ |\ \exists\,\tau_{k}\downarrow 0\ \ {\rm s.t.}\ \ {\rm dist}(x+\tau_{k}w,S)=o(\tau_{k})\big{\}},$
the regular normal cone to the set $S$ at the point $x$ is defined by
[…]
while the normal cone to $S$ in a given direction $d\in\mathbb{X}$ at $x$ is defined as
$\mathcal{N}_{S}(x;d):=\big{\{}v\in\mathbb{X}\ |\ \exists\,t_{k}\downarrow 0,\,d^{k}\to d,\,v^{k}\to v\ {\rm with}\ v^{k}\in\widehat{\mathcal{N}}_{S}(x\!+\!t_{k}d^{k})\big{\}}.$
For the mapping $R$, its graphical derivative mapping $DR(x)\!:\mathbb{X}\rightrightarrows\mathbb{X}$ is defined by $w\in DR(x)(u)$ iff $(u,w)\in\mathcal{T}_{{\rm gph}R}(x,R(x))$, and its (limiting) coderivative at $x$ in direction $(u,\xi)$ is defined by $w\in D^{*}R(x;(u,\xi))(v)$ iff $(w,-v)\in\mathcal{N}_{{\rm gph}R}(x;(u,\xi))$.
Proposition 2.2.
Fix any $\overline{x}\in\mathcal{S}^{*}$ with nonsingular $\nabla^{2}\!f(\overline{x})\!-\mathcal{I}$. Suppose that ${\rm prox}_{g}$ is directionally differentiable at $\overline{x}-\!\nabla\!f(\overline{x})$. If the directional pseudo-normality does not hold for system $R(x)=0$ at $\overline{x}$, then there exist $u\neq 0$ with $(u,-\nabla^{2}\!f(\overline{x})u)\in\mathcal{T}_{{\rm gph}\,\partial g}(\overline{x},\!-\!\nabla\!f(\overline{x}))$ and $v\neq 0$ with $(\nabla^{2}\!f(\overline{x})v,v)\in\!\mathcal{N}_{{\rm gph}\,\partial g}\big{(}(\overline{x},-\nabla\!f(\overline{x}));(u,-\nabla^{2}\!f(\overline{x})u)\big{)}$ such that $\langle(\nabla^{2}\!f(\overline{x})-\mathcal{I})v,u-{\rm prox}_{g}^{\prime}(\overline{x}-\!\nabla\!f(\overline{x});u-\!\nabla^{2}\!f(\overline{x})u)\rangle\geq 0$.
Proof 2.3.
Since the directional pseudo-normality does not hold for system $R(x)=0$ at $\overline{x}$, by invoking [5, Definition 4.1 (b)] with $P=R$ and $\Lambda=\{0\}$, there exist $0\neq(u,\xi)\in\widetilde{L}(\overline{x}):=\{(u,\xi)\in\mathbb{X}\times\mathbb{X}\ |\ \xi\in DR(\overline{x})(u)\cap\mathcal{T}_{\Lambda}(R(\overline{x}))\}$ and $\zeta\in\mathbb{X}\backslash\{0\}$ such that
$0\in D^{*}R(\overline{x};(u,\xi))(\zeta),\,\zeta\in\mathcal{N}_{\Lambda}(R(\overline{x});\xi)\ {\rm and}\ \exists\,(u^{k},s^{k},\zeta^{k})\to(u,R(\overline{x}),\zeta)$
${\rm and}\ t_{k}\downarrow 0\ {\rm such\ that}\ \zeta^{k}\in\widehat{\mathcal{N}}_{\Lambda}(s^{k})\ {\rm and}\ \langle\zeta,R(\overline{x}+t_{k}u^{k})\!-\!s^{k}\rangle>0.\qquad$
Note that $\mathcal{T}_{\Lambda}(R(\overline{x}))\!=\!\{0\},\,\mathcal{N}_{\Lambda}(R(\overline{x});\xi)\!=\!\mathbb{X}$ and $\widehat{\mathcal{N}}_{\Lambda}(0)=\mathbb{X}$. The above condition is equivalent to the existence of $u\neq 0$ with $0\in DR(\overline{x})(u)$ and $\zeta\neq 0$ such that
(7) $00\in D^{*}R(\overline{x};(u,\xi))(\zeta),\,\exists\,(u^{k},\zeta^{k})\to(u,\zeta)\ {\rm and}\ t_{k}\downarrow 0\ {\rm s.t.}\ \langle\zeta,R(\overline{x}+t_{k}u^{k})\rangle>$ 0 .
We next take a look at the tangent and normal cones to ${\rm gph}\,R$. By the expression of $R$ and ${\rm prox}_{g}=(\mathcal{I}+\partial g)^{-1}$, we have $R(x)-\!\nabla\!f(x)\in\partial g(x-\!R(x))$ for $x\in\mathbb{X}$, which implies that ${\rm gph}R=\Psi^{-1}({\rm gph}\,\partial g)$ with the mapping $\Psi\!:\mathbb{X}\times\mathbb{X}\to\mathbb{X}\times\mathbb{X}$ defined by
(8) $\Psi(x,y):=\left(\begin{matrix}x-y\\ y\!-\!\nabla\!f(x)\end{matrix}\right)\ \ {\rm for}\ (x,y)\in\mathbb{X}\times\mathbb{X}.$
Since $\nabla^{2}\!f(\overline{x})-\mathcal{I}$ is nonsingular, the Jacobian matrix $\Psi^{\prime}(\overline{x},0)$ is nonsingular, so there exists a neigborhood $V$ of $(\overline{x},0)$ such that for all $(x,y)\in V$, $\Psi^{\prime}(x,y)$ is nonsingular. By invoking [42, Exercise 6.7], for any $(x,y)\in{\rm gph}R\cap V$, it holds that
(9a) $\mathcal{T}_{{\rm gph}R}(x,y)=[\Psi^{\prime}(x,y)]^{-1}\mathcal{T}_{{\rm gph}\,\partial g}(\Psi(x,y)),$
(9b) $\widehat{\mathcal{N}}_{{\rm gph}R}(x,y)=\nabla\Psi(x,y)\,\widehat{\mathcal{N}}_{{\rm gph}\,\partial g}(\Psi(x,y))$
with $\nabla\Psi(x,y)\!=\![\Psi^{\prime}(x,y)]^{*}$, the adjoint of the Jacobian mapping. By (9a) and the definition of graphical derivative, $0\in\!DR(\overline{x})(u)$ iff $\Psi^{\prime}(\overline{x},0)(u,0)\in\mathcal{T}_{{\rm gph}\,\partial g}(\Psi(\overline{x},0))$, which is precisely $(u,-\nabla^{2}\!f(\overline{x})u)\in\mathcal{T}_{{\rm gph}\,\partial g}(\Psi(\overline{x},0))$ by equation (8). For the normal cone to ${\rm gph}R$ in the direction $(d_{x},d_{y})$ at $\overline{x}$, we claim that
(10) $\mathcal{N}_{{\rm gph}R}(\overline{x};(d_{x},d_{y}))\subset\nabla\Psi(\overline{x},0)\mathcal{N}_{{\rm gph}\,\partial g}\big{(}(\overline{x},-\nabla\!f(\overline{x}));(d_{x}\!-\!d_{y},d_{y}\!-\!\nabla^{2}\!f(\overline{x})d_{x})\big{)}.$
Indeed, pick any $(v_{x},v_{y})\!\in\!\mathcal{N}_{{\rm gph}R}(\overline{x};(d_{x},d_{y}))$. There are sequences $(d_{x}^{k},d_{y}^{k})\to(d_{x},d_{y})$, $t_{k}\downarrow 0,\,(v_{x}^{k},v_{y}^{k})\to(v_{x},v_{y})$ with $(v_{x}^{k},v_{y}^{k})\in\!\widehat{\mathcal{N}}_{{\rm gph}R}(\overline{x}\!+\!t_{k}d_{x}^{k},t_{k}d_{y}^{k})$ for each $k$. By equation (9b), $\widehat{\mathcal{N}}_{{\rm gph}R}(\overline{x}\!+\!t_{k}d_{x}^{k},t_{k}d_{y}^{k})=\nabla\Psi(\overline{x}\!+\!t_{k}d_{x}^{k},t_{k}d_{y}^{k})\widehat{\mathcal{N}}_{{\rm gph}\,\partial g}(\Psi(\overline{x}\!+t_{k}d_{x}^{k},t_{k}d_{y}^{k}))$ for all sufficiently large $k$. While from Assumption 1 (i), it holds that $\Psi(\overline{x}+t_{k}d_{x}^{k},t_{k}d_{y}^{k})=\big{(}\overline{x}+t_{k}(d_{x}^{k}-d_{y}^{k}),-\nabla\!f(\overline{x})+t_{k}(-\nabla^{2}\!f(\overline{x})d_{x}^{k}+d_{y}^{k}+o(t_{k})/t_{k})\big{)}$. Then the claimed inclusion (10) follows the definition of directional normal cone. Recall that $0\in D^{*}R(\overline{x};(u,0))(\zeta)$ iff $(0,-\zeta)\in\mathcal{N}_{{\rm gph}R}(\overline{x};(u,0))$. Together with inclusion (10), there exists $v$ such that $(\nabla^{2}\!f(\overline{x})v,v)\in\mathcal{N}_{{\rm gph}\,\partial g}\big{(}(\overline{x},-\nabla\!f(\overline{x}));(u,-\nabla^{2}\!f(\overline{x})u)\big{)}$ and $\zeta=(\nabla^{2}\!f(\overline{x})-\mathcal{I})v$. Note that ${\rm prox}_{g}$ is directionally differentiable in the Hadamard sense at $\overline{x}-\nabla\!f(\overline{x})$, so does the mapping $R$ at $\overline{x}$, which along with $R(\overline{x})=0$ implies that $R(\overline{x}+t_{k}u^{k})=t_{k}(R^{\prime}(\overline{x};u)+o(t_{k})/t_{k})$. Thus, condition (7) implies the desired conclusion.
Corollary 2.4.
Let $g(x)\!=\!\lambda\|x\|_{1}\ (\lambda\!>0)$ for $x\!\in\mathbb{R}^{n}$. Fix any $(\overline{x},\overline{y})\in{\rm gph}\,\partial g$. Write $J={\rm supp}(\overline{x})$ and $\overline{J}=\{1,2,\ldots,n\}\backslash{\rm supp}(\overline{x})$. The following results hold.
  • (i)
    if $(u,\xi)\in\mathcal{T}_{{\rm gph}\,\partial g}(\overline{x},\overline{y})$, for each $i\in\overline{J}$, we have $u_{i}>0,\xi_{i}=0$ or $u_{i}=0,\xi_{i}\leq 0$ if $\overline{y}_{i}=\lambda$; $u_{i}<0,\,\xi_{i}=0$ or $u_{i}=0,\,\xi_{i}\geq 0$ if $\overline{y}_{i}=-\lambda$; and $u_{i}=0$ if $|\overline{y}_{i}|<\lambda$.
  • (ii)
    $(u,\xi)\in\!\widehat{\mathcal{N}}_{{\rm gph}\,\partial g}(\overline{x},\overline{y})$ iff $u_{i}=0$ for each $i\in J$, and for each $i\in\overline{J}$, we have $u_{i}\in{\rm sign}(\overline{y}_{i})\,\mathbb{R}_{-},\xi_{i}\in{\rm sign}(\overline{y}_{i})\,\mathbb{R}_{+}$ if $|\overline{y}_{i}|=\lambda$; and $\xi_{i}=0$ if $|\overline{y}_{i}|<\lambda$.
  • (iii)
    The local error bound in (6) holds at $\overline{x}\in\mathcal{S}^{*}$ under the following (a) or (b):
    • (a)
      $|[\nabla\!f(\overline{x})]_{i}|<\lambda$ for $i\in\overline{J}$ and any $(u,-\nabla^{2}\!f(\overline{x})u)\in\mathcal{T}_{{\rm gph}\,\partial g}(\overline{x},-\nabla\!f(\overline{x}))$ satisfies $u_{J}=0$;
    • (b)
      $\nabla^{2}\!f(\overline{x})-\mathcal{I}$ and $[\nabla^{2}\!f(\overline{x})]_{JJ}$ are nonsingular and $|[\nabla\!f(\overline{x})]_{i}|<\lambda$ for $i\in\overline{J}$.
Proof 2.5.
By the separability of ${\rm gph}\,\partial g$ and [42, Proposition 6.41], an elementary calculation yields parts (i) and (ii). For part (iii), in accordance with the discussion after Definition 2.1, it suffices to argue that under condition (a) or (b), the directional pseudo-normality holds for system $R(x)=0$ at $\overline{x}$. By part (i), $u_{\overline{J}}=0$ if $|[\nabla\!f(\overline{x})]_{i}|<\lambda$ for $i\in\overline{J}$, so under (a) there is no $u\neq 0$ with $(u,-\nabla^{2}\!f(\overline{x})u)\in\mathcal{T}_{{\rm gph}\,\partial g}(\overline{x},-\nabla\!f(\overline{x}))$. Under (b), any $(\nabla^{2}\!f(\overline{x})v,v)\in\mathcal{N}_{{\rm gph}\,\partial g}\big{(}(\overline{x},-\nabla\!f(\overline{x}));(u,-\nabla^{2}\!f(\overline{x})u)\big{)}$ satisfies $v=0$, where $(u,-\nabla^{2}\!f(\overline{x})u)\in\mathcal{T}_{{\rm gph}\,\partial g}(\overline{x},-\nabla\!f(\overline{x}))$ with $u\neq 0$. Indeed, for such $(\nabla^{2}\!f(\overline{x})v,v)$, there exist $t_{k}\downarrow 0,\,(u^{k},\xi^{k})\to(u,-\nabla^{2}\!f(\overline{x})u)$ and $(s^{k},w^{k})\to(\nabla^{2}\!f(\overline{x})v,v)$ such that $(s^{k},w^{k})\in\widehat{\mathcal{N}}_{{\rm gph}\,\partial g}(\overline{x}\!+\!t_{k}u^{k},-\nabla\!f(\overline{x})+t_{k}\xi^{k})$. From the given condition $|[\nabla\!f(\overline{x})]_{i}|<\lambda$ for $i\in\overline{J}$, we have $|[\nabla\!f(\overline{x})-t_{k}\xi^{k}]_{i}|<\lambda$ for all $k$ large enough. By the expression of $\widehat{\mathcal{N}}_{{\rm gph}\,\partial g}$, we deduce that $s_{J}^{k}=0$ and $w_{\overline{J}}^{k}=0$. This means that $[\nabla^{2}\!f(\overline{x})v]_{J}=0$ and $v_{\overline{J}}=0$. Then, $[\nabla^{2}\!f(\overline{x})]_{JJ}v_{J}=0$, which by the nonsingularity of $[\nabla^{2}\!f(\overline{x})]_{JJ}$ implies that $v_{J}=0$ and $v=0$. Thus, the conclusion follows by Proposition 2.2.
Although the local superlinear convergence rate analysis of our algorithm depends on a local Hölderian error bound on the set $\mathcal{X}^{*}$ rather than $\mathcal{S}^{*}$, the above conditions are still instructive to examine the possibility of such an error bound to hold.
@@ 2.3 Kurdyka-Łojasiewicz property
By [3, Lemma 2.1], to show that a proper lsc function has the KL property (of exponent $\theta\in[0,1)$), it suffices to check if the property holds at its critical points. The discussions in [31, Section 2.2] show that the KL property of $h$ with exponent $\theta\in(0,1)$ has a close relation with the $\gamma$-subregularity of its subdifferential mapping $\partial h$ for $\gamma\in(0,1]$. As discussed in [3, Section 4], the KL property is ubiquitous and the functions definable in an o-minimal structure over the real field admit this property.
@@ 3 Inexact $q$-order regularized proximal Newton method
Input: $\epsilon\in(0,1),q\in[2,3],0<\!L_{m}<L_{\!M},\varrho\in(0,1),\sigma\in(0,1),\tau>1$ and $x^{0}\in{\rm dom}\,g$.
@@
  4. 4.
    If $r(x^{k+1})\leq\epsilon$ or $\|x^{k+1}\!-\!x^{k}\|\leq\epsilon$, stop; else go to step 1.
@@
(b) Algorithm 15 is well-defined by Lemma 3.2 later. When $x^{k+1}\!=x^{k}$ occurs, $x^{k}$ is a stationary point of (1) because now $x^{k}=y^{k}$ holds and the inexactness criterion (15) implies that $r(x^{k})=r_{k}(y^{k})=0$. Indeed, from (16), either $x^{k+1}=y^{k}$ or $x^{k+1}=y^{k}-v^{k}$. For the former, $x^{k}=y^{k}$ automatically holds; while for the latter, $x^{k}=y^{k}-v^{k}$, which by step (2b) implies that $\sigma(L_{k}/q)\|y^{k}\!-\!x^{k}\|^{q}\leq F(x^{k})\!-\!F(y^{k})\leq F(x^{k})\!-\!F(y^{k}\!-\!v^{k})=0$, so $x^{k}=y^{k}$ holds by $L_{k}\geq L_{m}>0$. This gives an explanation why $\|x^{k+1}\!-\!x^{k}\|\leq\epsilon$ is also adopted as a stopping condition of Algorithm 1.
@@
(d) Note that $v^{k}=R_{k,j_{k}}(y^{k})$ for each $k\in\mathbb{N}$. By the expression of $R_{k,j_{k}}$, we have
@@
    There exists $\overline{j}\in\mathbb{N}$ such that each subproblem (14) with $j\geq\overline{j}$ has a stationary point $\overline{y}^{k,j}$ satisfying $\Theta_{k,j}(\overline{y}^{k,j})<\Theta_{k,j}(x^{k})$.
@@
which means that $f_{k,j}(z)+g(z)\geq\vartheta_{k}(z)+\langle\widehat{\xi},z\rangle+(L_{k,j}/q)\|z-x^{k}\|^{q}+g(\widehat{x})-\langle\widehat{\xi},\widehat{x}\rangle$ for any $z\in\mathbb{X}$. Note that $L_{k,j}\to\infty$ as $j\to\infty$ due to $\tau>1$. Together with $q\in[2,3]$ and the expression of $\vartheta_{k}$, there must exist an index $\overline{j}\in\mathbb{N}$ such that for all $j\geq\overline{j}$,
$\vartheta_{k}(z)+\langle\widehat{\xi},z\rangle+(L_{k,j}/q)\|z-x^{k}\|^{q}+g(\widehat{x})-\langle\widehat{\xi},\widehat{x}\rangle\to\infty\ \ {\rm as}\ \|z\|\to\infty.$
That is, every subproblem (14) with $j\geq\overline{j}$ is coercive, so has an optimal solution. Let $\overline{y}^{k,j}$ be an optimal solution of (14) with $j\geq\overline{j}$. Clearly, $\overline{y}^{k,j}$ is a stationary point of subproblem (14) with $j\geq\overline{j}$ and $\Theta_{k,j}(\overline{y}^{k,j})\leq\Theta_{k,j}(x^{k})$. The rest only argues that $\Theta_{k,j}(\overline{y}^{k,j})\neq\Theta_{k,j}(x^{k})$. If not, then $x^{k}$ is an optimal solution of subproblem (14) with $j\geq\overline{j}$, and $0\in\nabla\!f_{k,j}(x^{k})+\partial g(x^{k})=\nabla\!f(x^{k})+\partial g(x^{k})$, a contradiction to $r(x^{k})\neq 0$.
(ii) Fix any $j\in\mathbb{N}$. Let $\overline{y}^{k,j}$ be a stationary point of (14) with $\Theta_{k}(\overline{y}^{k,j})<\Theta_{k}(x^{k})$. Clearly, $\overline{y}^{k,j}\neq x^{k}$. From $r_{k,j}(\overline{y}^{k,j})=0$, we have $r_{k,j}(\overline{y}^{k,j})-\varrho L_{k,j}\|\overline{y}^{k,j}\!-x^{k}\|^{q-1}<0$. Note that $r_{k,j}(\cdot)$ is continuous because ${\rm prox}_{g}$ is nonexpansive and $\nabla\!f_{k,j}$ is continuous. Then, for any $y$ sufficiently close to $\overline{y}^{k,j}$, $r_{k,j}(y)\!-\!\varrho L_{k,j}\|y\!-\!x^{k}\|^{q-1}<0$. The rest only argues that for any $y$ close enough to $\overline{y}^{k,j}$, $\Theta_{k}(y\!-\!R_{k,j}(y))\leq\Theta_{k}(x^{k})$. Indeed, from the continuity of $R_{k,j}$ and $R_{k,j}(\overline{y}^{k,j})=0$, we have $\big{(}y\!-\!R_{k,j}(y),R_{k,j}(y)\!-\!\nabla\!f_{k,j}(y)\big{)}\to\big{(}\overline{y}^{k,j},-\nabla\!f_{k,j}(\overline{y}^{k,j})\big{)}$ as $y\to\overline{y}^{k,j}$. In addition, by the expression of $R_{k,j}$, for any $z\in\mathbb{X}$, $R_{k,j}(z)-\nabla\!f_{k,j}(z)\in\partial g(z\!-\!R_{k,j}(z))$. Then, the subdifferential continuity of $g$ at $\overline{y}^{k,j}$ by [42, Example 13.30] implies that $g(y\!-\!R_{k,j}(y))\to g(\overline{y}^{k,j})$ as $y\to\overline{y}^{k,j}$. Thus, $\Theta_{k,j}(y-\!R_{k,j}(y))\to\Theta_{k,j}(\overline{y}^{k,j})$ as $y\to\overline{y}^{k,j}$. Together with $\Theta_{k,j}(\overline{y}^{k,j})<\Theta_{k,j}(x^{k})$, we have $\Theta_{k,j}(y-R_{k,j}(y))\leq\Theta_{k,j}(x^{k})$. The above arguments show that for any $y$ close enough to $\overline{y}^{k,j}$, the vector $y-R_{k,j}(y)$ satisfies condition (15).
@@
$F(x^{k})$ $>\vartheta_{k}(y^{k,j})+(L_{k,j}/q)\|y^{k,j}\!-\!x^{k}\|^{q}+\langle\widehat{\xi},y^{k,j}\!-\!\widehat{x}\rangle+g(\widehat{x}),$

### v4

@@ 2.2 Metric subregularity of residual mapping
(6) ${\rm dist}(x,\mathcal{S}^{*})\leq\kappa[r(x)]^{\gamma}\quad{\rm for\ all}\ x\in\mathbb{B}(\overline{x},\delta).$
This is the local Hölderian error bound of order $\gamma$ on $\mathcal{S}^{*}$ at $\overline{x}$, and is a little weaker than the one on $\mathcal{X}^{*}$ at $\overline{x}\in\mathcal{X}^{*}$. Some checkable conditions for the subregularity of $R$ are derived by using the results of [5]; see https://arxiv.org/abs/2311.06871v3.
@@ 2.3 Kurdyka-Łojasiewicz property
By [3, Lemma 2.1], to show that a proper lsc function has the KL property (of exponent $\theta\in[0,1)$), it suffices to check if the property holds at its critical points. The discussions in [30, Section 2.2] show that the KL property of $h$ with exponent $\theta\in(0,1)$ has a close relation with the $\gamma$-subregularity of its subdifferential mapping $\partial h$ for $\gamma\in(0,1]$. As discussed in [3, Section 4], the KL property is ubiquitous and the functions definable in an o-minimal structure over the real field admit this property.
@@ 3 Inexact $q$-order regularized proximal Newton method
Input: $q\in[2,3],0<\!L_{m}<L_{\!M},\varrho\in(0,1),\sigma\in(0,1),\tau>1$ and $x^{0}\in{\rm dom}\,g$.
@@
    and go to step 1.
@@
(b) Algorithm 11 is well-defined by Lemma 3.2 later. When $x^{k+1}\!=x^{k}$ occurs, $x^{k}$ is a stationary point of (1) because now $x^{k}=y^{k}$ holds and the inexactness criterion (11) implies that $r(x^{k})=r_{k}(y^{k})=0$. Indeed, from (12), either $x^{k+1}=y^{k}$ or $x^{k+1}=y^{k}-v^{k}$. For the former, $x^{k}=y^{k}$ automatically holds; while for the latter, $x^{k}=y^{k}-v^{k}$, which by step (2b) implies that $\sigma(L_{k}/q)\|y^{k}\!-\!x^{k}\|^{q}\leq F(x^{k})\!-\!F(y^{k})\leq F(x^{k})\!-\!F(y^{k}\!-\!v^{k})=0$, so $x^{k}=y^{k}$ holds by $L_{k}\geq L_{m}>0$. In view of this, we suggest that $r(x^{k+1})\leq\epsilon$ or $\|x^{k+1}\!-\!x^{k}\|\leq\epsilon$ for an accuracy tolerance $\epsilon\in(0,1)$ as the stopping criterion of Algorithm 1.
@@
Note that $v^{k}=R_{k,j_{k}}(y^{k})$ for each $k\in\mathbb{N}$. By the expression of $R_{k,j_{k}}$, we have
@@
    For each $j\in\mathbb{N}$, subproblem (10) has a stationary point $\overline{y}^{k,j}$ satisfying $\Theta_{k,j}(\overline{y}^{k,j})<\Theta_{k,j}(x^{k})$.
@@
which means that $f_{k,j}(z)+g(z)\geq\vartheta_{k}(z)+\langle\widehat{\xi},z\rangle+(L_{k,j}/q)\|z-x^{k}\|^{q}+g(\widehat{x})-\langle\widehat{\xi},\widehat{x}\rangle$ for any $z\in\mathbb{X}$. Note that $L_{k,0}>\max(0,-\lambda_{\min}(\nabla^{2}f(x^{k})))$ when $q=2$. Together with $q\in[2,3]$ and the expression of $\vartheta_{k}$, it follows that for each $j\in\mathbb{N}$, as $\|z\|\to\infty$,
$\vartheta_{k}(z)+\langle\widehat{\xi},z\rangle+(L_{k,j}/q)\|z-x^{k}\|^{q}+g(\widehat{x})-\langle\widehat{\xi},\widehat{x}\rangle\to\infty.$
That is, for each $j\in\mathbb{N}$, subproblem (10) is coercive, so has an optimal solution. Fix any $j\in\mathbb{N}$. Let $\overline{y}^{k,j}$ be an optimal solution of (10). Clearly, $\overline{y}^{k,j}$ is a stationary point of subproblem (10) and $\Theta_{k,j}(\overline{y}^{k,j})\leq\Theta_{k,j}(x^{k})$. The rest only argues that $\Theta_{k,j}(\overline{y}^{k,j})\neq\Theta_{k,j}(x^{k})$. If not, then $x^{k}$ is an optimal solution of subproblem (10), and $0\in\nabla\!f_{k,j}(x^{k})+\partial g(x^{k})=\nabla\!f(x^{k})+\partial g(x^{k})$, a contradiction to $r(x^{k})\neq 0$.
(ii) Fix any $j\in\mathbb{N}$. Let $\overline{y}^{k,j}$ be a stationary point of (10) with $\Theta_{k}(\overline{y}^{k,j})<\Theta_{k}(x^{k})$. Clearly, $\overline{y}^{k,j}\neq x^{k}$. From $r_{k,j}(\overline{y}^{k,j})=0$, we have $r_{k,j}(\overline{y}^{k,j})-\varrho L_{k,j}\|\overline{y}^{k,j}\!-x^{k}\|^{q-1}<0$. Note that $r_{k,j}(\cdot)$ is continuous because ${\rm prox}_{g}$ is nonexpansive and $\nabla\!f_{k,j}$ is continuous. Then, for any $y$ sufficiently close to $\overline{y}^{k,j}$, $r_{k,j}(y)\!-\!\varrho L_{k,j}\|y\!-\!x^{k}\|^{q-1}\leq 0$. The rest only argues that for any $y$ close enough to $\overline{y}^{k,j}$, $\Theta_{k}(y\!-\!R_{k,j}(y))\leq\Theta_{k}(x^{k})$. Indeed, from the continuity of $R_{k,j}$ and $R_{k,j}(\overline{y}^{k,j})=0$, we have $\big{(}y\!-\!R_{k,j}(y),R_{k,j}(y)\!-\!\nabla\!f_{k,j}(y)\big{)}\to\big{(}\overline{y}^{k,j},-\nabla\!f_{k,j}(\overline{y}^{k,j})\big{)}$ as $y\to\overline{y}^{k,j}$. In addition, by the expression of $R_{k,j}$, for any $z\in\mathbb{X}$, $R_{k,j}(z)-\nabla\!f_{k,j}(z)\in\partial g(z\!-\!R_{k,j}(z))$. Then, the subdifferential continuity of $g$ at $\overline{y}^{k,j}$ by [41, Example 13.30] implies that $g(y\!-\!R_{k,j}(y))\to g(\overline{y}^{k,j})$ as $y\to\overline{y}^{k,j}$. Thus, $\Theta_{k,j}(y-\!R_{k,j}(y))\to\Theta_{k,j}(\overline{y}^{k,j})$ as $y\to\overline{y}^{k,j}$. Together with $\Theta_{k,j}(\overline{y}^{k,j})<\Theta_{k,j}(x^{k})$, we have $\Theta_{k,j}(y-R_{k,j}(y))\leq\Theta_{k,j}(x^{k})$. The above arguments show that for any $y$ close enough to $\overline{y}^{k,j}$, the vector $y-R_{k,j}(y)$ satisfies condition (11).
@@
$F(x^{k})$ $\geq\vartheta_{k}(y^{k,j})+(L_{k,j}/q)\|y^{k,j}\!-\!x^{k}\|^{q}+\langle\widehat{\xi},y^{k,j}\!-\!\widehat{x}\rangle+g(\widehat{x}),$